the length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29cm, find the length of one of the equal legs.
A. 11cm
b. 4cm
b. 12cm
d. 5cm

Answer by Nate(3500)

base: b
one of the equal sides: 4b - 8
b + 4b - 8 + 4b - 8 = 29
9b - 16 = 29
9b = 45
b = 5
one of the equal sides: 4b - 8 = 20 - 8 = 12
(b)